# Drift2Matrix: Kernel-Induced Self Representation for Concept Drift Adaptation in Co-evolving Time Series

### v2

@@ H.8.4 Ablation Study on Kernel Functions
    $\bm{\mathcal{K}}(\mathbf{S},\mathbf{S})=\mathbf{S}^{\top}\mathbf{S}$ (26)
@@
    $\bm{\mathcal{K}}(\mathbf{S},\mathbf{S})=(\mathbf{S}^{\top}\mathbf{S}+c)^{d}$ (27)
@@
    $\bm{\mathcal{K}}(\mathbf{S},\mathbf{S}^{\prime})=\tanh(\xi\mathbf{S}^{\top}\mathbf{S}^{\prime}+c)$ (28)
@@
As shown in Table 8, the Drift2Matrix model with the Gaussian kernel achieves the best performance across all datasets, indicating its effectiveness in capturing the complex nonlinear relationships in time series data. The ablation study highlights the Gaussian kernel’s ability to adaptively handle various types of data distributions, making it a versatile choice for time series analysis. However, the choice of the kernel function may depend on the specific characteristics of the dataset, and therefore, a careful consideration of kernel properties is necessary for optimal model performance.
Table 8: Ablation study of Drift2Matrix with different kernel functions, in terms of RMSE, for the nine datasets
Dataset Drift2Matrix (Linear) Drift2Matrix (Polynomial) Drift2Matrix (Sigmoid) Drift2Matrix
@@
$\bm{\mathcal{K}}=\sum_{m=1}^{M}\beta_{m}\bm{\mathcal{K}}_{m},\quad\text{subject to}\quad\sum_{m=1}^{M}\beta_{m}=1$
where $\bm{\mathcal{K}}_{m}$ represents individual kernels and $\beta_{m}$ are the non-negative weights that sum to 1. This formulation allows MKL to find an optimal combination of kernels.

### v3

@@ G.8.4 Ablation Study on Kernel Functions
    $\boldsymbol{\mathcal{K}}(\mathbf{S},\mathbf{S})=\mathbf{S}^{\top}\mathbf{S}$ (26)
@@
    $\boldsymbol{\mathcal{K}}(\mathbf{S},\mathbf{S})=(\mathbf{S}^{\top}\mathbf{S}+c)^{d}$ (27)
@@
    $\boldsymbol{\mathcal{K}}(\mathbf{S},\mathbf{S}^{\prime})=\tanh(\xi\mathbf{S}^{\top}\mathbf{S}^{\prime}+c)$ (28)
@@
As shown in Table 8, the CORAL model with the Gaussian kernel achieves the best performance across all datasets, indicating its effectiveness in capturing the complex nonlinear relationships in time series data. The ablation study highlights the Gaussian kernel’s ability to adaptively handle various types of data distributions, making it a versatile choice for time series analysis. However, the choice of the kernel function may depend on the specific characteristics of the dataset, and therefore, a careful consideration of kernel properties is necessary for optimal model performance.
Table 8: Ablation study of CORAL with different kernel functions, in terms of RMSE, for the nine datasets
Dataset CORAL (Linear) CORAL (Polynomial) CORAL (Sigmoid) CORAL
@@
$\boldsymbol{\mathcal{K}}=\sum_{m=1}^{M}\beta_{m}\boldsymbol{\mathcal{K}}_{m},\quad\text{subject to}\quad\sum_{m=1}^{M}\beta_{m}=1$
where $\boldsymbol{\mathcal{K}}_{m}$ represents individual kernels and $\beta_{m}$ are the non-negative weights that sum to 1. This formulation allows MKL to find an optimal combination of kernels.